# DifFlow3D: Toward Robust Uncertainty-Aware Scene Flow Estimation with Iterative Diffusion-Based Refinement

### v2

@@ 4.3 Plug-and-Play on Previous Works
[…]
@@ 4.4 Ablation Study
Uncertainty. As illustrated in Table 4 (b), without our proposed uncertainty, there is $5.0\%$ higher EPE3D. Our uncertainty ground truth is designed as a binary value (0 or 1). Here, we also conduct the experiment by substituting it with a continuous value ranging from 0 to 1. However, worse flow estimation results can be seen. This is because binary uncertainty can better distinguish accurate and inaccurate estimation and further lead the network to optimize inaccurate flows. To better guide the flow regression, we jointly estimate uncertainty and scene flow as in Section 3.3. If we predict uncertainty and scene flow by two separate fully connected layers from cost volume, there is a $21.4\%$ EPE3D increase. This is because uncertainty can not tightly control the flow due to the indirect connection and even mislead the flow learning with the wrong direction.
Condition signals. Conditional information is extremely significant in guiding the flow residual generation in our diffusion-based refinement module. Here, we remove each of our designed condition signals to validate their importance as in Table 4 (c). Without cost volume as guidance, our model has obviously worse estimation (3.4 times larger EPE3D), since cost volume precisely represents the per-point correlation between two frames. Coarse flow embedding can also lead the flow generation because the coarse matching in the low-resolution layer has a more global understanding of cross-frame correlation. Thus, if we remove it, there is a $3.7\%$ accuracy drop in terms of Acc3DS. Geometry feature has low-level 3D structural information in dynamic scenes. $7.1\%$ higher EPE3D can be seen without the geometry feature as guidance.
@@
Ours (with GRU-based $M_{\theta}$ [5]) 0.0140 0.9776 0.9933 0.0479
@@
Ours (full, with all three conditions) 0.0140 0.9776 0.9933 0.0479
@@ 5 Discussion
We discuss the motivation and reasons why our proposed diffusion-based refinement module can improve the scene flow estimation accuracy in this section.
Why diffusion? As illustrated in Fig. 1, the introduction of diffusion model can improve the robustness to some challenging cases, such as dynamics, noisy inputs, repetitive patterns, etc. Here, we further show two quantitative comparisons by adding random noise in Fig. 6. Previous works all witness a dramatically increasing EPE3D, but our method has relatively small fluctuation in terms of estimation accuracy.
How uncertainty works? Our uncertainty represents the reliability and confidence level of per-point predicted scene flow, which can progressively constrain the flow residual generation range within diffusion model. As in Fig. 7, there are large uncertainty intervals in terms of all points at initial stages of the training. This means that our network has learned less prior to distinguish correct or incorrect flow matching. After training for 100 epochs as in Fig. 7 (b), all of uncertainty intervals narrow obviously because our network has better estimation ability. Furthermore, uncertainty intervals also vary in terms of different objects. As shown in Fig. 7 (b), flows of the dynamic car are relatively difficult for the model to learn due to its inconsistent motion. Thus, it can be seen a relatively larger uncertainty interval, while other static objects have almost no uncertainty estimates.
@@ 6 Conclusion
In this paper, we innovatively propose a diffusion-based scene flow refinement network with awareness of estimation uncertainty. Multi-scale diffusion refinement is adopted to generate fine-grained dense flow residuals. To improve the robustness of our estimation, we also introduce a per-point uncertainty jointly generated with scene flow. Extensive experiments demonstrate the superiority and generalization ability of our DifFlow3D. Notably, our diffusion-based refinement can serve as a plug-and-play module into previous works and shed new light on future works.

### v3

@@ 4.3 Plug-and-Play on Previous Works
It is extremely encouraging that our diffusion-based refinement can serve as a plug-and-play module, improving the estimation accuracy of several recent SOTA works [44, 4, 5]. In Table 3, we replace their refinement modules with our diffusion-based one. 3DFlow [44] first trains their model on a quarter of the data, then finetunes results on the complete dataset. For the comprehensive comparison, we also show experiment results on both quarter and complete training sets. As shown in Table 3, our method respectively has 6.3% and 13.9% EPE3D reduction on the quarter and complete training set of FT3Ds dataset. Similarly, introducing our diffusion refinement strategy into the GRU in BPF [4] and MSBRN [5] can also significantly improve their accuracy. Also, they innovate the cross-frame association with specific designs. Instead, our diffusion-based refinement is more generic in terms of different network designs or challenging cases. As in Fig. 5, estimations for dynamic cars or bushes with complex patterns are more accurate by adding our Diffusion-based Scene Flow Refinement (DSFR).
@@ 4.4 Ablation Study
Uncertainty. As illustrated in Table 4 (b), without our proposed uncertainty, there is $28.9\%$ higher EPE3D. Our uncertainty ground truth is designed as a binary value. We also conduct the experiment by substituting it with a continuous value ranging from 0 to 1. However, worse estimation results can be seen. To better guide the flow regression, we jointly estimate uncertainty and scene flow as in Section 3.3. If we predict uncertainty and scene flow by two separate fully connected layers from cost volume, there is a significant EPE3D increase. This is because uncertainty can not tightly control the flow due to the indirect connection and even mislead the flow learning in the wrong direction.
Condition signals. Conditional information is extremely significant in guiding the flow residual refinement module. We remove each of our designed condition signals to validate their importance as in Table 4 (c). Without cost volume as guidance, our model has obviously worse estimation (4.5 times larger EPE3D), since cost volume precisely represents the per-point correlation between two frames. Coarse flow embedding can also lead the flow generation because the coarse matching in the low-resolution layer has a more global understanding of cross-frame correlation. Thus, if we remove it, there is a significant accuracy drop. The geometry feature has low-level 3D structural information in dynamic scenes. $31.6\%$ higher EPE3D can be seen without the geometry feature as guidance.
@@
Ours (with GRU-based $M_{\theta}$ [5]) 0.0114 0.9836 0.9949 0.0350
@@
Ours (full, with all three conditions) 0.0114 0.9836 0.9949 0.0350
@@
4.5 Runtime Comparison
We compare our DifFlow3D with previous methods on efficiency. All results are evaluated on a single RTX 3090 GPU. Table 5 shows that DifFlow3D has highly competitive efficiency. Notice that our baseline here is based on MSBRN [5], where we combine its flow initialization with our diffusion-based refinement. Compared with MSBRN, our method has only a slightly more computational overhead ($3.3\%$), but much higher estimation accuracy ($24.0\%$ and $29.1\%$ on FT3Ds and KITTIs as in Table 1).
@@ 5 Discussion
We discuss the motivations and reasons why our proposed diffusion-based refinement module works.
Why diffusion? As illustrated in Fig. 1, the introduction of diffusion model can improve the robustness for some challenging cases. We further show two quantitative comparisons by adding random noise in Fig. 6. Previous works all witness a dramatically increasing EPE3D, but our method has a relatively small fluctuation in accuracy.
How uncertainty works? Our uncertainty represents the reliability and confidence level of per-point predicted scene flow, which can progressively constrain the flow residual generation range. As in Fig. 7, there are large uncertainty intervals for all points at the initial stages of the training. After training for 100 epochs as in Fig. 7 (b), all uncertainty intervals are narrow obviously because our network has better estimation ability. Furthermore, uncertainty intervals also vary in terms of different objects. As shown in Fig. 7 (b), there are relatively larger uncertainty intervals for the dynamic car, which are relatively difficult to learn due to its inconsistent motion. But other static objects have almost no uncertainty estimations.
@@ 6 Conclusion
In this paper, we innovatively propose a diffusion-based scene flow refinement network with uncertainty awareness. Multi-scale diffusion refinement is adopted to generate fine-grained dense flow residuals. To improve the robustness of our estimation, we also introduce a per-point uncertainty jointly generated with scene flow. Extensive experiments demonstrate the superiority and generalization ability of our DifFlow3D. Notably, our diffusion-based refinement can serve as a plug-and-play module into previous works and shed new light on future works.
Acknowledgement. This work was supported in part by the Natural Science Foundation of China under Grant 62225309, 62073222, U21A20480, and 62361166632.